OpenStudy (anonymous):

|10-3x|+5=2 (Absolute Zero)

OpenStudy (anonymous):

[ 10 - 3x ] + 5 = 2 10-3x+5=2 -3x+15=2 -3x=2-15 -3x=-13 x=13/3

OpenStudy (anonymous):

Can I ask what grade level You are?

OpenStudy (anonymous):

Thats not it. It has to do with Absolute Zero

OpenStudy (anonymous):

Oh ok let me fix it